OpenStudy (anonymous):

WILL GIVE A MEDAL HELP (a) A liter of paint covers an area of 9 square meters. It takes 3 liters of paint to paint a rectangular wall that is 7 meters high. How wide is the wall ? (b) A liter of paint covers an area of A square meters. It takes v liters of paint to paint a rectangular wall that is h meters high. How wide is the wall ? (a) width of wall = _________ meters (b) width of wall = ________ meters

jimthompson5910 (jim_thompson5910):

" A liter of paint covers an area of 9 square meters. It takes 3 liters of paint to paint a rectangular wall " so what is the area of the entire wall?

OpenStudy (anonymous):

If one liter of paint is 9 square meters, and it takes 3 liters of paint the wall, you will have to multiply 9 by 3, which ends up with 27.

jimthompson5910 (jim_thompson5910):

correct, the wall is 27 square meters

jimthompson5910 (jim_thompson5910):

we have this wall 

jimthompson5910 (jim_thompson5910):

x is the width of the wall

jimthompson5910 (jim_thompson5910):

how can we find x?

OpenStudy (anonymous):

27/7

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

cool ok

jimthompson5910 (jim_thompson5910):

A liter of paint covers an area of A square meters. It takes v liters of paint to paint a rectangular wall that is h meters high. How wide is the wall ?

jimthompson5910 (jim_thompson5910):

This is the same idea, but now just using unknown variables instead of numbers

OpenStudy (anonymous):

wait thats it for 1A???

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

part a)

OpenStudy (anonymous):

okay

jimthompson5910 (jim_thompson5910):

"A liter of paint covers an area of A square meters. It takes v liters of paint to paint a rectangular wall that is h meters high" the wall is A*v square meters the height is h 

jimthompson5910 (jim_thompson5910):

what is x?

OpenStudy (anonymous):

so we would solve for x then?

OpenStudy (anonymous):

x=h/A

jimthompson5910 (jim_thompson5910):

Area = Length * Width Av = h*x x = Av/h

OpenStudy (anonymous):

oh rightt _._ I see my mistake

OpenStudy (anonymous):

thank you, I now understand the problem!

jimthompson5910 (jim_thompson5910):

I'm glad it's clicking